# ConKeD: Multiview contrastive descriptor learning for keypoint-based retinal image registration

### v1

@@ 1 Introduction
Image registration is the process of spatially aligning a pair of images. In this process, one image is used as the reference, the fixed image, while the second image, the moving image, is spatially transformed to match the first one. In the medical imaging domain, these images are usually from the same subject and they differ in aspects such as the point of view from which they were taken or the time at which they were captured. In medicine, image registration is particularly important due to its multiple applications on clinical practice Viergever et al. (2016). Image registration enables the simultaneous analysis of multiple images which, in turn, allows clinicians to draw more informed conclusions Hajnal et al. (2001). For instance, medical image registration enables longitudinal studies by aligning images taken at different times, which is useful to monitor disease progression or remission Narasimha-Iyer et al. (2007). Similarly, automatic registration methods are fundamental for computer-aided diagnosis systems Morita (2019); Yanase and Triantaphyllou (2019) as these cannot rely on manual registration since it is a complex and time-intensive procedure.
Retinal Image Registration (RIR) is a highly relevant task as the eyes are the only organs in the human body that allow non-invasive in vivo observation of the blood vessels and neuronal tissue Forrester et al. (2020). In particular, there is a high interest in the registration of color fundus (CF) images due to the widespread use of this image modality in clinical practice Hernandez-Matas et al. (2021). CF images are cost effective due to the relatively low price of CF cameras combined with their efficacy in the diagnosis of numerous diseases Ho et al. (2018). For instance, CF images allow to diagnose conditions such as Diabetic Retinopathy, Glaucoma or Age-Related Macular Degeneration Salmon (2020). However, CF images also present particular characteristics that complicate the registration. For instance, their photographic nature can cause several imperfections due to incorrect placement of the device, incorrect capture by the clinician or movements by the patient. Some of these imperfections are color or illuminance variation, changes in focus, halos, etc. Moreover, certain diseases can drastically alter the retinal appearance. These morphological changes complicate the registration process in the presence of pathology. The combination of all of these characteristics makes the CF registration process a challenging task.
In broad terms, automatic registration approaches can be classified according to their algorithmic basis as either classical approaches or deep learning approaches. Furthermore, registration methods can also be divided into three main groups: Feature Based Registration (FBR), which uses keypoints and descriptors; Intensity Based Registration (IBR), which uses the intensity values of the image; and Direct Parameter Regression (DPR), which predicts a deformation field or matrix directly from the input images Rivas-Villar et al. ; Rivas-Villar et al. (2023a).
@@
IBR approaches work by maximizing a similarity metric between the intensity values of the images to be registered across transformation parameter space. These methods are commonly iterative processes optimizing the the similarity metrics with approaches such as gradient descent. There are many suitable similarity metrics for this task, depending on the application field, image modalities, etc. Pluim et al. (2000); Balakrishnan et al. (2018). Recently, deep learning methods have been applied in this registration paradigm by learning new similarity metrics with Convolutional Neural Networks (CNNs) Cheng et al. (2018).
@@
DPR methods allow to directly predict the transformation parameters or the deformation field that aligns the images Haskins et al. (2020); Balakrishnan et al. (2019). To achieve this, a deep neural network is trained such that its output is used to transform the moving image. In this case, the training objective is to maximize the similarity between the transformed moving image and the fixed image or, equivalently, between two simplified representations of these images (such as e.g. blood vessel segmentation maps for CF images Benvenuto et al. (2022)). The key difference between IBR and DPR is how they use similarity metrics. IBR explicitly maximizes a similarity metric while DPR uses the similarity metric to guide the learning of neural network. This network learns to create suitable transformations directly from the input data, only relying on the similarity metric as guidance during its training step.
@@
Finally, there are hybrid methods joining several of these registration paradigms. In this case, FBR methods usually provide an initial alignment which is then refined by other IBR or DPR methods that are capable of deformable registration thus improving the results Álvaro S. Hervella et al. (2018); Rivas-Villar et al. (2023b). However, while these approaches might be successful in terms of registration results they are, generally, not desirable as they greatly increase the computational costs as well the number of hyperparameters which are needed to tune the methods. Furthermore, if the initial FBR registration does not provide an accurate enough registration the subsequent refinement step will fail to improve the performance.
@@
Currently, most state-of-the-art medical registration methodologies are based on IBR or DPR methods Balakrishnan et al. (2019); de Vos et al. (2019); Xiao et al. (2021). However, RIR tasks, and particularly CF images, have certain characteristics that limit the applicability of these methods. In particular, the patterns that are useful for the registration of CF images (i.e. blood vessel bifurcations and crossovers, optic disc, etc.) are relatively small and are scattered over a homogeneous background. Furthermore, CF images can present large displacements and low degree of overlapping between the images in each pair. Additionally, there can be multiple intra-pair differences produced by the photographic nature of the capture process (i.e., light halos, blurriness due to suboptimal focus, etc.). These changes are specifically detrimental to IBR methods as creating a robust similarity metric is challenging under these conditions. Similarly, as the transformations on the retina are almost always rigid, deformable registration methods like DPR using deformation fields are not suitable as they pose the risk of overfitting these fields creating unrealistic transformations. Therefore, commonly used methods in other medical image modalities such as Balakrishnan et al. (2019); de Vos et al. (2019) are unsuitable for color fundus registration.
@@
Traditionally, the CF registration field has been dominated by classical FBR methods Hernandez-Matas et al. (2020); Motta et al. (2019). Some of these methods, such as REMPE and VOTUS, still obtain the best results in most of the categories of the well known FIRE benchmark dataset Hernandez-Matas et al. (2020); Motta et al. (2019). REMPE Hernandez-Matas et al. (2020) combines generic points (SIFT) and domain-specific keypoints (blood vessel bifurcations). Then, using RANSAC (Random Sample Consensus) Fischler and Bolles (1981) and Particle Swarm Optimization, the transformation is computed according to an ellipsoidal model of the eye that has been created specifically for this task. Differently, VOTUS Motta et al. (2019) creates graphs for the arterio-venous tree. Then, the graphs from the corresponding images are matched using a novel algorithm (Vessel Optimal Transform) that relies on classical image features. The transformation matrix is computed using DeSAC (Deterministic Sample And Consensus).
@@
Recently, some novel FBR deep learning based methods have been able to compete with and even outperform the classical approaches in some cases Rivas-Villar et al. (2022, ); Liu et al. (2022), whereas previous DPR deep learning methods could not Zou et al. (2020). Some of these methods are based on domain-specific keypoints Rivas-Villar et al. (2022); Liu et al. (2022) while others employ generic keypoint detectors Rivas-Villar et al. .
@@
The work of Rivas-Villar et. al. Rivas-Villar et al. (2022) was the first FBR deep learning method for CF image registration. It uses a CNN to detect blood vessel crossovers and bifurcations which are then used to infer a transformation using RANSAC without the need for descriptor computation. However, the lack of descriptors limits the complexity of the transformation as calculating all the possible combinations of keypoints with RANSAC by brute-force increases in computational cost with the degrees of freedom in the geometric transformation. SuperRetina Liu et al. (2022) adapted SuperPoint DeTone et al. (2018), a natural image registration method, to RIR. They propose to train this network using keypoints labeled by an expert manually or by a classical approach, known as PBO Oinonen et al. (2010). To perform the matching of keypoints, they use descriptors that are learnt using a triplet loss approach. Their method requires extensive ad-hoc image pre-processing, firstly taking only the green channel from the RGB images, normalizing it, applying CLAHE contrast enhancement Pizer et al. (1987), and finally using gamma adjustment. These pre-processing steps simplify the task but require multiple parameters that may need to be tuned on a per-dataset or per-device basis. After this, the keypoints and descriptors are jointly trained starting from a set of labeled keypoints which are progressively expanded using the more confident detections of the network outside of the ground truth. In order to complete the registration pipeline, SuperRetina uses a double inference step. This means that, the registration is firstly done normally and, after transforming the moving image, the process is repeated again to improve the results at the expense of more computation cost. Similarly, in Rivas-Villar et al. R2D2 Revaud et al. (2019) is adapted for the retinal image registration task. R2D2 is a natural image registration method capable of jointly learning keypoints and descriptors without ground truth data, making this method fully unsupervised. Unlike SuperRetina, this method does not require pre-processing or double inference. However, it obtains worse results. In this method the descriptors are learnt using a multi-negative loss, known as AP loss.
@@
Overall, using domain-specific keypoints, such as blood vessel crossovers and bifurcations, involves the use of supervised learning and manually labeled data Rivas-Villar et al. (2022); Liu et al. (2022), whereas methods which use generic keypoint detectors do not require labeled data Rivas-Villar et al. . However, due to the lack of precise and highly distinctive keypoints, generic detectors usually carry an increased computational cost. In terms of descriptor learning, deep learning methods use contrastive learning, a technique based on comparing samples. For a given keypoint, which is used as reference (anchor keypoint), there are positive samples (the same point in another view of the same image) and negative samples (any other point in the same or another view). The main difference among methods is how they use the different samples and how many samples they are able to process on each contrastive learning step. Currently, in natural image registration as well as in RIR, most approaches use triplet loss Liu et al. (2022); DeTone et al. (2018) which exploits a single positive and negative sample per anchor. Moreover, in RIR, there are some approaches that employ multi-negative single-positive methods Rivas-Villar et al. or do not use descriptors at all Rivas-Villar et al. (2022). In this regard, none of the state-of-the-art methods use novel multi-positive multi-negative approaches for image registration, neither in natural images nor in medical images. Multiple positive contrastive learning has been successfully used in other tasks, such as image classification, where it has demonstrated improved performance over just multi-negative approaches Khosla et al. (2020). These methods employ multiviewed batches to create multiple positives for each anchor, which in turn, allows to leverage more information from the input images. However, no multi-positive multi-negative framework has been developed for keypoint-based registration despite the success and benefits that this self-supervised approach demonstrated in other areas He et al. (2020); Chen et al. (2020).
@@
  • Our proposal achieves results comparable to those in the state-of-the-art while avoiding pre-processing, requiring less training samples, and also requiring less detected keypoints.
@@ 2.1 Methodology
The proposed methodology for CF registration is described in Fig. 1. This methodology is based on ConKeD, our novel framework for learning descriptors for keypoint-based image registration. These keypoints are detected using a state-of-the-art neural network Rivas-Villar et al. (2022). Meanwhile, another network is used to compute the descriptor of these keypoints. This network is trained with a multiviewed batch containing one original image and $N$ views or augmentations of the original image. Therefore, each keypoint (i.e. anchor) will have $N$ positive samples and all the rest of detected keypoints act as negatives including the points in the own image. Using all of these samples in conjunction with a suitable loss the keypoint description is trained, learning to match the positive samples and differentiate the negatives. At inference time, the keypoints are detected, described, matched and, finally, RANSAC is used to infer a projective transform (homography) to align both images.
@@ 2.1.1 Keypoint detection
In order to detect the domain-specific keypoints, we train a CNN to generate heatmaps depicting all the possible keypoints in the input image Álvaro S. Hervella et al. (2020a). In this case, the use of heatmaps instead of binary maps is beneficial due to the heavy imbalance between the positive class (the keypoint pixels) and the negative class (the background pixels)Álvaro S. Hervella et al. (2020a). The produced heatmaps have maximum values at the location of the detected keypoints and progressively decreasing values in the neighboring pixels. The heatmaps increase the information available to the network during training, improving the learning process Álvaro S. Hervella et al. (2020a).
@@
Finally, during inference, in order to extract the keypoints, we compute the heatmaps using this network. From the heatmaps we can obtain the discrete locations of the keypoints by using a local maxima filter and an intensity threshold Álvaro S. Hervella et al. (2020a).
@@ 2.1.2 Keypoint description
ConKeD follows a novel multi-positive multi-negative contrastive learning approach in order to learn descriptors for image registration. Typically, deep learning-based FBR methods learn descriptors using triplet loss approaches. These approaches compare each given point against both a positive sample and a negative sample. Some recent methods also explored multi-negative approaches which, instead of selecting a particular negative sample, take advantage of all of them. In contrast, our method leverages multiple samples for both positive and negative examples. This avoids the need for explicit mining and drastically increases the amount of contrastive comparisons, facilitating the network training.
The fundamental idea behind our proposal is to compare the representations (i.e, descriptors) of different keypoints. In particular, the key concept is to pull together the anchor (a particular point used as reference) and all the positive samples in the descriptor space while pushing apart the anchor from all the different negative samples.
@@
In order to compare the descriptors, we use cosine similarity as the metric. Similarly to other works Chen et al. (2020); van den Oord et al. (2018); Khosla et al. (2020), we apply L2 normalization to the output of the neural network which in combination with the dot product used in the losses described below is equivalent to cosine similarity. Thus, the cosine similarity between two descriptors can be directly computed as the dot product. Moreover, this normalization helps contrastive losses to perform intrinsic hard positive/negative mining Khosla et al. (2020).
As training loss for ConKeD we propose two different alternatives: SupConLoss Khosla et al. (2020) and MP-InfoNCE van den Oord et al. (2018); He et al. (2020); Chen et al. (2020).
@@
SupCon Loss Khosla et al. (2020) is a loss function specifically designed for multi-positive multi-negative contrastive learning in the context of supervised image classification. This loss function can be directly applied to our proposal. For our framework, which uses a multiviewed batch of $N+1$ images, each of them containing $K$ keypoints, let $S$ be the set of all the keypoints such that its cardinality is $|S|=(N+1)K$. In this case, $s\in S$ is the index of an arbitrary anchor keypoint taken from $S$. Let $P(s)$ be the set of the corresponding positive samples for $s$ (i.e. other views of the anchor keypoint) and $p\in P$ be one of these positive samples. Similarly, $A(s)=S\setminus{s}$ is the set of all the keypoints except the anchor (i.e. both the positive and the negative samples for $s$).Therefore, the SupCon loss function is defined as:
$\mathcal{L}_{sup}=\sum_{s\in S}{\frac{-1}{|P(s)|}\sum_{p\in P(s)}{\log\frac{exp(z_{s}\cdot z_{p}/\tau)}{\sum\limits_{a\in A(s)}{exp(z_{s}\cdot z_{a}/\tau)}}},}$ (1)
where we represent the network output for any keypoint as $z$, and we use the symbol $\cdot$ to denote the dot product. Finally, $\tau\in R$ is the scalar temperature parameter.
@@
InfoNCE van den Oord et al. (2018), also known as NT-Xent, is a loss function commonly used in many metric learning methods van den Oord et al. (2018); He et al. (2020); Chen et al. (2020). However, it is only suitable for the paradigm of a single positive and many negatives. Therefore we propose to modify it so that it can be applied in our framework. We define our version of InfoNCE, MP-InfoNCE as follows.
$\mathcal{L}_{NCE}=\sum_{i\in I}{\frac{1}{|I^{\prime}|}\sum_{x\in X_{i}}{\frac{1}{|X_{i}|}\sum_{\begin{subarray}{c}j\in I\\ j>i\end{subarray}}{-\log\frac{exp(z_{x}\cdot z_{p_{j}(x)})/\tau)}{\sum\limits_{c\in C_{i,j}(x)}{exp(z_{x}\cdot z_{c})/\tau}}}}},$ (2)
where $I$ is the set of images which compose our multiviewed batch such that its cardinality is $|I|=(N+1)$ $I^{\prime}$ denotes the subsets of two elements of $I$ such that its cardinality is $|I^{\prime}|={I\choose 2}$. We define $i\in I$ and $j\in I$ such that $j>i$ are the indexes of two different images from $I$ forming a pair in $I^{\prime}$. Let $X_{i}$ and $X_{j}$ be the sets of all keypoints in images $i$ and $j$. In this case, $x$ is the index of an arbitrary anchor keypoint taken from $X_{i}$ and $p_{j}(x)$ is its corresponding positive sample from $X_{j}$. Additionally, $C_{i,j}(x)=(X_{i}\cup X_{j})\setminus x$ is the set of all the keypoints in images $i$ and $j$ except for the anchor keypoint (i.e. all the negatives and the positive sample).

### v2

@@ 1 Introduction
Image registration is the process of spatially aligning a pair of images. In this process, one image is used as the reference, the fixed image, while the second image, the moving image, is spatially transformed to match the first one. These images are usually from the same subject and they differ in their point of view or in the instant they were captured in.  In medicine, image registration is particularly important due to its multiple applications on clinical practice [1]. Image registration enables the simultaneous analysis of multiple images which, in turn, allows clinicians to draw more informed conclusions [2]. Additionally, medical image registration enables longitudinal studies which are useful to monitor disease progression or remission [3].
Retinal Image Registration (RIR) is a highly relevant task as the eyes are the only organs in the human body that allow non-invasive in vivo observation of the blood vessels and neuronal tissue [4]. In particular, there is a high interest in the registration of color fundus (CF) images due to the widespread use of this image modality in clinical practice [5]. CF images are cost effective due to the relatively low price of CF cameras combined with their efficacy in the diagnosis of numerous diseases [6, 7]. However, CF images also present particular characteristics that complicate the registration. For instance, their photographic nature can cause several imperfections due to incorrect placement of the device, incorrect capture by the clinician or movements by the patient. Some of these imperfections are color or illuminance variation, changes in focus, halos, etc. Moreover, certain diseases can drastically alter the retinal appearance. These morphological changes complicate the registration process in the presence of pathology. The combination of all of these characteristics makes the CF registration process a challenging task.
In broad terms, automatic registration approaches can be classified according to their algorithmic basis as either classical approaches or deep learning approaches. Furthermore, registration methods can also be divided into three main groups: Feature Based Registration (FBR), which uses keypoints and descriptors; Intensity Based Registration (IBR), which uses the intensity values of the image; and Direct Parameter Regression (DPR), which predicts a deformation field or matrix directly from the input images [8, 9].
@@
IBR approaches work by iteratively maximizing a similarity metric between the intensity values of the images to be registered across transformation parameter space. There are multiple suitable similarity metrics [10, 11], including deep learning-based ones [12].
@@
DPR methods allow direct transformation prediction or deformation field prediction which, in turn, allows image alignment [13, 14]. To achieve this, a deep neural network is trained such that its output is used to transform the moving image. In this case, the training objective is to maximize the similarity between the transformed moving image and the fixed image or, equivalently, between two simplified representations of these images (such as e.g. blood vessel segmentation maps for CF images [15]). The key difference between IBR and DPR is how they use similarity metrics. IBR explicitly maximizes a similarity metric while DPR uses the similarity metric to guide the learning of neural network. This network learns to create suitable transformations directly from the input data, only relying on the similarity metric as guidance during its training step.
@@
Finally, there are hybrid methods joining several of these registration paradigms. In this case, FBR methods usually provide an initial alignment which is then refined by other IBR or DPR methods that are capable of deformable registration thus improving the results [16, 17]. However, while these approaches might be successful in terms of results, they are not desirable since they increase the complexity and the computational cost and may introduce extra hyperparameters.
@@
Currently, most state-of-the-art medical registration methodologies are based on IBR or DPR methods [14, 18, 19]. However, RIR tasks, and particularly CF images, have certain characteristics that limit the applicability of these methods. In particular, the patterns that are useful for the registration of CF images (i.e. blood vessel bifurcations and crossovers, optic disc, etc.) are relatively small and are scattered over a homogeneous background. Furthermore, CF images can present large displacements and low degree of overlapping between the images in each pair. Additionally, there can be multiple intra-pair differences produced by the photographic nature of the capture process (i.e. light halos, blurriness due to suboptimal focus, etc.). These changes are specifically detrimental to IBR methods as creating a robust similarity metric is challenging under these conditions. Similarly, as the transformations on the retina are almost always rigid, deformable registration methods like DPR using deformation fields are not suitable as they pose the risk of overfitting these fields creating unrealistic transformations. Therefore, commonly used methods in other medical image modalities such as [14, 18] are unsuitable for color fundus registration.
@@
Traditionally, the CF registration field has been dominated by classical FBR methods [20, 21]. Some of these methods, such as REMPE and VOTUS, still obtain the best results in most of the categories of the well-known FIRE benchmark dataset [20, 21]. REMPE [20] combines generic points (SIFT) and domain-specific keypoints (blood vessel bifurcations). Then, using RANSAC (Random Sample Consensus) [22] and Particle Swarm Optimization, the transformation is computed according to an ellipsoidal model of the eye that has been created specifically for this task. Differently, VOTUS [21] creates graphs for the arterio-venous tree. Then, the graphs from the corresponding images are matched using a novel algorithm (Vessel Optimal Transform) that relies on classical image features. The transformation matrix is computed using DeSAC (Deterministic Sample And Consensus).
@@
Recently, some novel FBR deep learning based methods have been able to compete with and even outperform the classical approaches in some cases [23, 8, 24], whereas previous DPR deep learning methods could not [25]. Some of these methods are based on domain-specific keypoints [23, 24] while others employ generic keypoint detectors [8].
@@
The work of Rivas-Villar et. al. [23] was the first FBR deep learning method for CF image registration. It uses a CNN to detect blood vessel crossovers and bifurcations which are then used to infer a transformation using RANSAC without the need for descriptor computation. However, the lack of descriptors limits the complexity of the transformation as calculating all the possible combinations of keypoints with RANSAC by brute-force increases in computational cost with the degrees of freedom in the geometric transformation. SuperRetina [24] adapted SuperPoint [26], a natural image registration method, to RIR. They propose to train this network using keypoints labeled by an expert manually or by a classical approach, known as PBO [27]. To perform the matching of keypoints, they use descriptors that are learnt using a triplet loss approach. Their method requires extensive ad-hoc image pre-processing, firstly taking only the green channel from the RGB images, normalizing it, applying CLAHE contrast enhancement [28], and finally using gamma adjustment. These pre-processing steps simplify the task but require multiple parameters that may need to be tuned on a per-dataset or per-device basis. After this, the keypoints and descriptors are jointly trained starting from a set of labeled keypoints which are progressively expanded using the more confident detections of the network outside of the ground truth. In order to complete the registration pipeline, SuperRetina uses a double inference step. This means that, the registration is firstly done normally and, after transforming the moving image, the process is repeated again to improve the results at the expense of more computation cost. Similarly, in [8] R2D2 [29] is adapted for the retinal image registration task. R2D2 is a natural image registration method capable of jointly learning keypoints and descriptors without ground truth data, making this method fully unsupervised. Unlike SuperRetina, this method does not require pre-processing or double inference. However, it obtains worse results. In this method the descriptors are learnt using a multi-negative loss, known as AP loss.
@@
Overall, using domain-specific keypoints, such as blood vessel crossovers and bifurcations, involves the use of supervised learning and manually labeled data [23, 24], whereas methods which use generic keypoint detectors do not require labeled data [8]. However, due to the lack of precise and highly distinctive keypoints, generic detectors usually carry an increased computational cost. In terms of descriptor learning, deep learning methods use contrastive learning, a technique based on comparing samples. For a given keypoint, which is used as reference (anchor keypoint), there are positive samples (the same point in another view of the same image) and negative samples (any other point in the same or another view). The main difference among methods is how they use the different samples and how many samples they are able to process on each contrastive learning step. Currently, in natural image registration as well as in RIR, most approaches use triplet loss [24, 26] which exploits a single positive and negative sample per anchor. Moreover, in RIR, there are some approaches that employ multi-negative single-positive methods [8] or do not use descriptors at all [23]. In this regard, none of the state-of-the-art methods use novel multi-positive multi-negative approaches for image registration, neither in natural images nor in medical images. Multiple positive contrastive learning has been successfully used in other tasks, such as image classification, where it has demonstrated improved performance over just multi-negative approaches [30]. These methods employ multiviewed batches to create multiple positives for each anchor, which in turn, allows to leverage more information from the input images. However, no multi-positive multi-negative framework has been developed for keypoint-based registration despite the success and benefits that this self-supervised approach demonstrated in other areas [31, 32].
@@
  • Our proposal achieves results comparable to those in the state-of-the-art while avoiding image pre-processing, requiring less training samples, and also requiring less detected keypoints.
@@ 2.1 Methodology
The proposed methodology for CF registration is described in Fig. 1. This methodology is based on ConKeD, our novel framework for learning descriptors for keypoint-based image registration. These keypoints are detected using a state-of-the-art neural network [23]. Meanwhile, another network is used to compute the descriptors of these keypoints. The descriptor network is trained with a multiviewed batch containing one original image and $N$ views or augmentations of the original image. Therefore, each keypoint (i.e. anchor) will have $N$ positive samples while all the rest of detected keypoints act as negatives, including the points in the own image. These samples coupled with a suitable loss allow the keypoint descriptor to be trained, learning to match the positive samples and differentiate the negatives. At inference time, the keypoints are detected, described, matched and, finally, RANSAC is used to infer a projective transform (homography) to align both images.
@@ 2.1.1 Keypoint detection
In order to detect these keypoints, we train a CNN to generate heatmaps depicting all the possible keypoints in the input image [33]. Using heatmaps instead of binary maps is beneficial due to the heavy imbalance between the positive class (the keypoint pixels) and the negative class (the background pixels)[33]. The produced heatmaps have maximum values at the location of the detected keypoints and progressively decreasing values in the neighboring pixels. The heatmaps increase the information available to the network during training, improving the learning process [33].
@@
Finally, during inference, in order to extract the keypoints, we compute the heatmaps using this network. From the heatmaps we can obtain the discrete locations of the keypoints by using a local maxima filter and an intensity threshold [33].
@@ 2.1.2 Keypoint description
ConKeD follows a novel multi-positive multi-negative contrastive learning approach in order to learn descriptors for image registration. Typically, deep learning-based FBR methods learn descriptors using triplet loss approaches. These approaches compare each given point against both a positive and a negative sample. Some recent methods also explored multi-negative approaches which, instead of selecting a particular negative sample, use of all of them. In contrast, our method leverages multiple samples for both positive and negative examples. This avoids the need for explicit mining and increases the amount of contrastive comparisons, facilitating training.
The fundamental idea behind our proposal is to compare the representations (i.e. descriptors) of different keypoints. In particular, the key concept is to pull together the anchor (a particular point used as reference) and all the positive samples in the descriptor space while pushing apart the anchor from all the different negative samples.
@@
In order to compare the descriptors, we use cosine similarity as the metric. Similarly to other works [32, 34, 30], we apply L2 normalization to the output of the neural network which in combination with the dot product used in the losses described below is equivalent to cosine similarity. Thus, the cosine similarity between two descriptors can be directly computed as the dot product. Moreover, this normalization helps contrastive losses to perform intrinsic hard positive/negative mining [30].
As training loss for ConKeD we propose two different alternatives: SupCon Loss [30] and MP-InfoNCE [34, 31, 32]. An schematic comparison between both losses can be seen in Figure 3.
Figure 3: Schematic comparison between SupCon Loss and MP-InfoNCE, considering N+1 views, where N is the number of augmented versions of the original (1) image, and K keypoints per view, each of which matches with one keypoint on each other view (i.e. N positives per keypoint/sample).
@@
SupCon Loss [30] is a loss function specifically designed for multi-positive multi-negative contrastive learning in the context of supervised image classification. This loss function can be directly applied to our proposal. For our framework, which uses a multiviewed batch of $N+1$ images, each of them containing $K$ keypoints, $k$ is the unique index of an arbitrary keypoint such that $k\in[1,K]$. Therefore, the SupCon Loss function is defined as:
$\mathcal{L}_{sup}=\frac{1}{N}\sum_{i=1}^{N+1}{\sum\limits_{\begin{subarray}{c}j=1,\\ j\neq i\end{subarray}}^{N+1}}\sum_{k=1}^{K}-\log\frac{exp(z_{ik}\cdot z_{jk}/\tau)}{\sum\limits_{\begin{subarray}{c}c=1,\\ c\neq k\end{subarray}}^{K}{exp(z_{ik}\cdot z_{ic}/\tau)}+\sum\limits_{\begin{subarray}{c}l=1,\\ l\neq i\end{subarray}}^{N+1}{\sum\limits_{c=1}^{K}{exp(z_{ik}\cdot z_{lc}/\tau)}}},$ (1)
where we represent the network output for any keypoint as $z$. This way, $z_{ik}$, $z_{jk}$, $z_{lk}$ represent the output of network for keypoint $k$ in the views of index $i$, $j$ and $l$, respectively. We use the symbol $\cdot$ to denote the dot product. Finally, $\tau\in R$ is the scalar temperature parameter.
@@
InfoNCE [34], also known as NT-Xent, is a loss function commonly used in many metric learning methods [34, 31, 32]. However, it is only suitable for the paradigm of a single positive and many negatives. Therefore we propose to modify it so that it can be applied in our framework. We define our version of InfoNCE, MP-InfoNCE as follows.
$\mathcal{L}_{NCE}=\frac{1}{{N+1\choose 2}K}\sum_{i=1}^{N+1}\sum_{\begin{subarray}{c}j=1,\\ j>i\end{subarray}}^{N+1}\sum_{k=1}^{K}{-\log\frac{exp(z_{ik}\cdot z_{jk}/\tau)}{\sum\limits_{\begin{subarray}{c}c=1,\\ c\neq k\end{subarray}}^{K}{exp(z_{ik}\cdot z_{ic}/\tau)}+\sum\limits_{c=1}^{K}{exp(z_{ik}\cdot z_{jc}/\tau)}}},$ (2)
where $N+1$ is the number of views in the our multiviewed batch ($1$ original image, $N$ augmented views). Since we compute the loss pairwise, the total number of combinations amounts to (${N+1\choose 2}$. $i$ and $j$, such that $j>i$, are the indexes of two different images from the multiviewed batch forming a pair. $k$ represents the unique index of a keypoint such that $k\in[1,K]$. This way, $z_{ik}$ represents the output of network for keypoint $k$ in the view of index $i$. Likewise, $z_{jk}$ represents the output of network for keypoint $k$ in the view of index $j$.
Overall, the main differences between SupCon Loss and MP-Info-NCE lie in the denominator of the innermost summation term. The denominator in SupCon Loss contains the matching similarities among all keypoints across all views, involving as many positives as augmented views N. Conversely, the denominator of MP-Info-NCE only contrasts against the matching similarities within the views involved in the numerator, and only involving one positive from one augmented view. Note that for N=1, i.e. 2 images, both losses would be equivalent.